OpenStudy (reneeshumpert):

2y = x + 3 5y = x - 7 What is the solution set of the given system? {(-1/3, -4/3)} {(1/3, 4/3)} {(-29/3, -10/3)} {(29/3, 10/3)} can someone explain how I do questions like this? I'm not good at all at fractions.

OpenStudy (pinklion23):

hi ill help

OpenStudy (reneeshumpert):

Okay , I'm just confused on how I solve the equations

OpenStudy (pinklion23):

umm what grade math is this

OpenStudy (reneeshumpert):

10th

OpenStudy (pinklion23):

oh sorry i cant help i thought i could but im in a lower grade sorry again i will call some one @iGreen

OpenStudy (reneeshumpert):

Thanks for trying at least

OpenStudy (pinklion23):

thx for the medal lol

OpenStudy (anonymous):

There are many ways how to teach fractions, and the success or failure of the student to understand fractions will depend on the teaching method used. The methods outlined below to teach fractions have been proven to be easy for the student to grasp, and easy for every math teacher to implement. It's been said that if a student understands fractions, then they can understand any mathematics concept. It is then very important for every math teacher to know how to teach fractions in the most approachable way possible. How to Teach Fractions Tip #1 - Explain every fraction in terms of pizzas to make fractions familiar. When deciding on a method of how to teaching fractions, we need to use fractional analogies that the student will immediately recognize. Thus enters the pizza as the perfect instrument needed to teach the concept of the fraction. How to Teach Fractions Tip #2 - Explain the numerator and the denominator until every student understands. Students will never understand how to add or subtract a fraction until they truly understand the concept of a fraction. Use fractions that every student understands. For example, take the fraction: Take the time to explain that this fraction represents a pizza that has been cut into two pieces, but you only have one of the pieces. The numerator of the fraction is the top number and the denominator of the fraction is the bottom number. Always keep in mind that using the best methods of how to teach fractions involves moving very slowly, giving multiple examples involving fractions until every student is comfortable. Use familiar examples involving fractions such as 1/4, 1/3, and 3/4 to reinforce the concept of what a fraction is. How to Teach Fractions Tip #3 - Draw lots of pictures of fractions. Children are visual learners and the best technique of how to teach fractions involves drawing pictures of fractions. For example, take the fraction: This fraction can be best described by drawing a pizza cut into four pieces, but you only have three of the pieces such as in this picture of the fraction:

OpenStudy (reneeshumpert):

no problem

OpenStudy (camerondoherty):

We have two equations We can solve for x and y by using substitution 2y = x + 3 5y = x - 7 Fist, we're going to use the first equation and isolate x to do that, we subtract 3 from both sides to get:\[x = 2y - 3 \]Now we plug this for x into the second equation and solve for y So:\[5y = (2y-3)-7 \]Can you solve that?